Opened 11 years ago

Closed 9 years ago

#213 closed flight processing (fixed)

VOC_10, flight day 314/2008, Chile

Reported by: benj Owned by: benj
Priority: alpha 4 high Milestone: 2008 data processing completion
Component: Archiving Keywords:
Cc: Other processors:

Description

Data location: /data/larsenb1/scratch/arsf/vocals/vocals-Q/voc10-31408 (symlink in ~arsf/arsf_data/2008/flight_data/vocals/VOC_10-2008_314_Chile)

Data arrived from ARSF via SATA disk Q on 12th December 2008.

Scientific details: VOCALS cloud campaign

Priority:

PI: G. Vaughan

Sensors:

  • Eagle
  • Hawk

Quicklook

TBD.

Change History (14)

comment:1 Changed 11 years ago by jatt

Have renamed Eagle files so all start with VNIR prefix.

comment:2 Changed 10 years ago by jatt

The first part of flight line 10 for both eagle and hawk contains numerous bends and has been processed to level 1 only.

comment:3 Changed 10 years ago by jatt

Flight lines 1 and 16 are CM 69.
Flight lines 2 - 15 are CM 75.

The pitch, roll and heading values used for this flight are:-

Sensor: Eagle

Flight Line Pitch Roll Heading
01-16 -0.2 -0.45 -0.5

Sensor: Hawk

Flight Line Pitch Roll Heading
01 -0.238 0.116 0.786
02 -0.245 0.105 0.436
03 -0.284 0.126 0.739
04 -0.253 0.121 0.779
05a -0.183 0.13 0.722
05b -0.183 0.13 0.722
06 -0.263 0.135 0.832
07 -0.214 0.176 0.798
08 -0.217 0.169 0.8
09a -0.26 0.271 0.601
09b -0.26 0.271 0.601
10 -0.217 0.316 0.881
11 -0.267 0.319 0.604
12a -0.294 0.337 0.813
12b -0.294 0.337 0.813
13 -0.255 0.356 0.654
14 -0.288 0.104 0.767
15 -0.317 0.093 0.661
16 -0.296 0.324 0.842

comment:4 Changed 10 years ago by jatt

  • Owner changed from jatt to emca

Delivery directory created at:-

~arsf/arsf_data/2008/flight_data/vocals/VOC_10-2008_314_Chile/delivery/20090827/VOC10

comment:5 Changed 10 years ago by jatt

Hawk line 16 - still has dark frames in the level 1 file.

Amended the autodarkstartline in the hawk header file, new value is 18535.

Re-processed and new files copied to the delivery directory for check.

comment:6 Changed 10 years ago by mark1

checked the hawk files, have reprocessed some to remove darkframes from level 1 files. All hawk files look good now.

comment:7 Changed 10 years ago by emca

Delivery check complete. Eagle lines 1, 10, 9b and 16 have been reprocessed to remove dark frames.

Ready for delivery.

comment:8 Changed 10 years ago by mark1

All data delivered 17th September 2009

comment:9 Changed 10 years ago by benj

  • Component changed from Processing: general to Archiving
  • Owner changed from emca to benj

Requested archiving by DTG 24th Sep. 2009

comment:10 Changed 10 years ago by benj

Archived by NEODC 5/10/2009, ready to go to NEODC

comment:11 Changed 10 years ago by benj

Er... whoops, meant archived by DTG 5/10/2009

comment:12 Changed 9 years ago by benj

Raw tarball copied to USB disk 45 12/02/2010 ready for transfer to NEODC. Delivery tarball yet to copy.

comment:13 Changed 9 years ago by benj

Delivery tarball copied to USB disk 47 16/02/2010 ready for transfer to NEODC.

comment:14 Changed 9 years ago by benj

  • Resolution set to fixed
  • Status changed from new to closed

Disk posted to NEODC 18/02/2010

Note: See TracTickets for help on using tickets.